i have 4 one cent stamps 3 five cent stamps and 3 twenty five cent stamps. how many different combinations can be made with these stamps? Answer is 89 but whats the equation? Nick? 8th grade math.
Answer by ikleyn(52794)

.
i have 4 one cent stamps 3 five cent stamps and 3 twenty five cent stamps. how many different combinations can be made
with these stamps? Answer is 89 but whats the equation? Nick? 8th grade math.
~~~~~~~~~~~~~~~~~~~~

Let me reformulate the problem in more clear way.

"I have 4 one-cent coins; 3 five-cents coins; and 3 twenty-five cent coins.
How many different sums can I combine using these coins?"

1.  You can get the sums of 1 cent, 2 cents, 3 cents and 4 cents 
    by combining one 1-cent coin, two 1-cent coins, three 1-cent coins and four 1-cent coins, respectively.
    4 different sums.

2.  You can get three sums of 5 cents, 10 cents, and 15 cents by combining . . . five-cent coins.
    3 different sums.

3.  You can get three sums of 25 cents, 50 cents, and 75 cents by combining . . . twenty-five-cent coins.
    3 different sums.

4.  4*3 = 12 different sums by combining 4 one-cent coins with 3 five-cent coins.

5.  4*3 = 12 different sums by combining 4 one-cent coins with 3 twenty-five-cent coins.

6.  3*3 = 9 different sums by combining 3 five-cent coins with 3 twenty-five-cent coins.

7.  4*3*3 = 36 different sums by combining 3 five-cent coins with 3 five-cent coins and 3 twenty-five-cent coins.

8. Now add all these combinations:

   4 + 3 + 3 + 12 + 12 + 9 + 36 = 79.

9. It is my answer.